# Do you monitor CI Practices? I don’t know. You tell me: A case study

### v1

@@ 4.5. RQ4: Are developers interested in using the tool?
Figure 3 shows the access data of the three organizations analyzed in the case study. The red bars represent tool access on the days when interviews were conducted. The blue bars represent access on days without interviews. The red crossed-out bars represent days when interviews took place, but no access was recorded in the tool.
@@
Figure 3. (A) Number of Access to the monitoring tool of Organization A during the case study. (B) Number of Access to the monitoring tool of Organization B during the case study. (C) Number of Access to the monitoring tool of Organization C during the case study.
@@
In Organization A, during the case study, we observed a total of 43 accesses to the tool dashboard of 2 different developers. Notably, the primary access points coincided with interview dates, as indicated by red bars on the timeline. These interview-centric peaks suggest heightened activity during those periods, potentially linked to collaborative discussions or focused on project evaluations. However, it is noteworthy that many accesses occurred outside the interview dates, indicating ongoing engagement and utilization of the tool beyond the case study evaluation periods. This broader pattern of access underscores the continuous and diverse interactions with the monitoring tool, emphasizing its relevance throughout the case study period.
@@
In Organization B, we observed 31 accesses to the tool dashboard by two different participants. Although 3 participants participated in the case study, one did not access the tool. We notice many accesses outside the interview dates (blue bars), indicating interest in monitoring CI practices. There is a sharp decline in access to the monitoring tool after December 22, 2023. This decline can be attributed to two factors: (i) the project going on a holiday break; and (ii) one of the participants taking vacation in the last two weeks of the case study. Another consideration is that the project was in a phase with few changes. Combined with receiving the values of CI practices via email, this caused a decrease in interest in accessing the tool interface, as explained by participant P3: “I received the email and the issue of code coverage appeared, as I already knew about code coverage, I did not pay much attention to the tool because I thought that perhaps it would be more of a priority to look at the issue of the pipeline, … , for the tool to return the collection of the coverage metric correctly.”
@@
In the last project, Organization C, there was the more reduced access to the monitoring tool, with only 7 accesses to the tool dashboard by 1 participant. This could be justified by the company’s cultural resistance to CI adoption. Additionally, the project had a holiday break at the end of the year, which is reflected in the access curve dropping and rising again when Organization Activities resumed. Another observation is that only one participant had access to the monitoring tool during the case study. This could potentially hinder the shift in the company’s culture regarding CI.
@@
The registration of accesses to the dashboard screen allows us to conclude that projects A and B, which had defined themselves as having an INTERMEDIATE CI maturity level, showed greater interest in accessing the tool. Organization C, which has a BASIC CI maturity, exhibited a reduced interest in using the tool.
4.6. RQ5: What was the evolution of CI practices during the use of the tool?
@@
Figure 4. (A) Organization A CI Practices Evolution during the case study. (B) Organization B CI Practices Evolution during the case study. (C) Organization C CI Practices Evolution during the case study.
@@
In Organization A, since the initiation of our monitoring efforts, all CI practices have demonstrated considerable improvement. Notably, the “Time to Fix a Broken Build” experienced a significant reduction, indicative of enhanced efficiency in addressing issues promptly. As a consequence, the “Build Health” has increased significantly. The improvement in test coverage is equally relevant, and interviews revealed that monitoring practices played a crucial role in driving this positive trend. Participant P1 said: “I think the evaluation is good for us to have more detailed information, so we can make changes to the project, where necessary. We already knew about the coverage problem from the beginning, but it became very evident. With this information, we began to act to resolve this problem.”. The “Commits per Day” practice dropped in the middle of the case study, followed by a strong recovery. We associate it to the departure of a team member and, subsequently, the entry of new members into the development team. The same event of change in the development team may have influenced the “Comments per Change” practice, which increased during the case study.
@@
Organization B presented a very unstable CI environment due to the complex integration environment with other applications, reflected in the CI practices. The code coverage mostly remained zero despite the project having automated tests. The build time fluctuated greatly between normal weeks and weeks with very high peaks (over 80 hours long). Participant P3 reflected on the issues with build time and coverage: “Yes, we have brought attention to these aspects in terms of build duration and code coverage. We are trying to understand this aspect in meetings, until this morning, the project managers asked me why the tests are taking so long”. This was a problem they were unable to resolve during the case study. This instability is also reflected in the high variation of “Build Health” and “Time to Fix a Broken Build” practices, as Participant P2 declares: “Last week we had some problems with infrastructure, we ended up having to do some configurations and such. Because of it, the build health metric started to have a lower value”. The “Commits per Day” tended to decrease in the last 4 weeks of the case study compared to the first 4 weeks. This trend can be attributed to the conclusion of the development cycle, during which the project shifted its focus toward documentation. As reinforced by participant P4: “… as everyone is focused on documentation, there shouldn’t be any more commits, except something from the README, something very specific like that.”
@@
In Organization C, the project initially demonstrated good values for CI practices, which was surprising considering the self-declared BASIC level of CI in the initial survey. However, there was a significant degradation in all metrics during the break for the end-of-year holidays (the red dotted circle). Despite this, the project experienced a quick recovery after the holiday period for most of the practices, such as: “Build Health”, “Build Activity” and “Time to Fix a Broken Build”. The CI monitoring helped with this recovery after the holiday break as Participant 5 said: “There was also a period at the end of last year of the build being generated in a broken way, and as soon as we came back, one of the first things we did was fix it… We had already observed this in Gitlab, but the monitoring tool also contributed, in a way, complemented this need to fix it”
@@
Organization A experimented with a notable enhancement in CI practices throughout the case study, particularly in coverage practices where reports indicate that monitoring played a crucial role in identifying and addressing areas that needed improvement. On the other hand, Organization B faced challenges due to the instability of its CI environment, impairing significant improvements in the project’s CI levels. Finally, in Organization C, the break for the end-of-year holiday has caused a temporary reduction in CI practices, which were quickly improved with the help of monitoring.
5. Discussion
In this section, we discuss the implications of our study.
@@ 5. Discussion
During the conduction of the case study, it was observed that the automation of the build process and the execution of automated tests are the primary factors considered when defining the CI maturity level of projects. However, when CI monitoring was introduced, developers emphasized the monitoring process. Monitoring highlighted problems, served as a reminder, and motivated the team to enhance the CI process. They have a clear perception that this improvement could have an impact on project quality.
@@
Combining CI practices into a single tool gives an overview of the CI process. It would be important for tools like GitLab to incorporate monitoring of CI practices. This integration would help in monitoring CI maturity. The call for tools like GitLab to incorporate monitoring of CI practices reflects a desire for more seamless and integrated solutions. Direct integration of CI monitoring into popular developer platforms, such as GitLab, streamlines the workflow for development teams. It supports efficient code collaboration and provides real-time insights into the health and effectiveness of CI processes. This integration aligns with the broader industry movement towards DevOps practices, where the collaboration between development and operations is tightly integrated.
Implications for researchers:
@@
This case study showed that there are still gaps in monitoring development processes. Even though CI is already a widespread practice, few studies seek to evaluate its correct application. There is a field of study exploring new CI practices not evaluated in this study or even broader practices such as DevOps.
Implications for practitioners:
@@
This study demonstrated that developers highly desire the monitoring of CI practices and derive significant benefits from its adoption. Therefore, commercial CI services like GitHub, GitLab, or Jenkins would greatly benefit from incorporating dashboards to monitor development practices such as CI practices. This addition would add substantial value to these tools and contribute to improving the quality of projects utilizing them.
6. Threats to Validity
In this section, we discuss the limitations of our study and how we attempted to mitigate them. Based on the work of (Elazhary et al., 2021), we utilized the framework outlined by (Roller and Lavrakas, 2015) to organize this section. This framework examines the sources of variability and possible bias in qualitative methods.
6.1. Data Gathering
This section focuses on discussing the accuracy and reliability of the collected data.
@@ 6.1. Data Gathering
Scope: While our case study has provided valuable insights into Continuous Integration practices, the scope of our study was limited to three software development organizations, in our immediate geographic area, so we may have physical access and permission to install the monitoring suite and collect data. The projects were represented by 5 participants. As a result, the generalization of our findings may be constrained to the project’s context, and the participants can not express a whole project understanding of CI. We tried to select participants who had knowledge about the project and the project’s CI infrastructure, as a way to mitigate this problem.
@@
Period of Case Study: The period in which the study was conducted - 8 weeks - may not be enough to have a complete effect on any action taken by the team after becoming aware of the CI project maturity. Furthermore, as we were immersed in the project’s environment, no developer was dedicated exclusively to the case study, continuing with their daily tasks, so they may not have had time to implement changes in the analyzed project, even though they were alerted by the monitoring of CI practices.
@@
The case study was carried out from November 2023 until February 2024. This period coincides with the holiday season. Therefore, some projects were at the end of the development cycle, with lower demand, final tests or project documentation were carried out, which had a slight impact on some metrics during a certain project period. A break in interviews was held between 2023-12-22 and 2024-01-19 to mitigate this problem.
@@
Changes in the Project’s Team: Organization A had changes in its development team during the case study, which may have influenced the evolution of CI practice values in RQ5. An intern left the Organization A at the beginning of the case study and three new interns entered at the end of the case study (these interns did not participate in the case study). In Organization B, one of the participants had a vacation during the last 2 weeks of the study, which may have influenced the data in RQ4 and RQ5.
6.2. Data Analysis
This section is dedicated to examining the accuracy of the analysis performed and the interpretations drawn from the study.
@@ 6.2. Data Analysis
7. Conclusion
In this paper, we conducted a case study across three projects to investigate the benefits of monitoring CI practices in real-world scenarios. The monitoring of CI practices remains a research gap and is often overlooked by the industry. Our main objective was to address this gap in practical settings. To achieve this, we developed a monitoring suite tool deployed in the projects’ environments, providing study participants with the ability to monitor CI practices with minimal intrusion.
The findings from our study revealed that monitoring CI practices plays a crucial role in the software development lifecycle. It highlights potential issues within the CI process and provides an overview of the CI states, serving as a constant motivator and reminder for developers to enhance their CI processes. Our study indicates developers’ strong desire to integrate CI monitoring dashboards with popular CI services tools, such as GitLab, GitHub, and Jenkins, suggesting that seamless integration could further optimize and facilitate the development workflow.
In future work, it is highly recommended to expand the scope of CI monitoring practices to incorporate additional DevOps practices, such as metrics from the DORA (DevOps Research and Assessment)  [^3] https://dora.dev/. Extending the monitoring framework to include deployment and infrastructure metrics offers an extended perspective on the software delivery lifecycle. Including these broader metrics aligns with industry best practices and ensures a more robust evaluation of DevOps practices’ overall efficiency and effectiveness within a project.
Acknowledgements.
This work is partially supported by INES (www.ines.org.br), CNPq grant 465614/2014-0, CAPES grant 88887.136410/2017-00, and
FACEPE grants APQ-0399-1.03/17 and PRONEX APQ/0388-1.03/14.

### v2

@@ IV-D RQ4: Are developers interested in using our metrics suite tool?
Figure 2 shows the tool access data of the three organizations analyzed in the case study. The red bars represent tool access on the days when interviews were conducted. The blue bars represent access on days without interviews. The red crossed-out bars represent days when interviews took place, but no access was recorded in the tool.
@@
Figure 2: (A) Number of Access to the monitoring tool of Organization A during the case study. (B) Number of Access to the monitoring tool of Organization B during the case study. (C) Number of Access to the monitoring tool of Organization C during the case study.
@@
In Organization A, during the case study, we observed a total of 43 accesses to the tool dashboard of 2 different developers. In particular, the primary access points coincided with interview dates, as indicated by the red bars on the timeline. These interview-centric peaks suggest increased activity during these periods, potentially linked to collaborative discussions or focused on project evaluations. However, it should be noted that many accesses occurred outside the interview dates, indicating ongoing engagement and usage of the tool beyond the case study evaluation periods. This broader pattern of access underscores the continuous and diverse interactions with the monitoring tool, emphasizing its relevance throughout the case study period.
@@
In Organization B, we observed 31 accesses to the tool dashboard by two different participants. Although three participants took part in the case study, one did not use the tool. We noticed many accesses outside the interview dates (blue bars), indicating interest in monitoring CI practices. There was a sharp decline in access to the monitoring tool after December 22, 2023. This decline can be attributed to two factors: (i) the project going on a holiday break, and (ii) one participant taking vacation during the last two weeks of the case study. Another consideration is that the project was in a phase with few changes. Combined with receiving CI practice values via email, this caused a decrease in interest in accessing the tool interface, as explained by Participant P3: “I received the email and the issue of code coverage appeared. Since I was already aware of code coverage, I did not pay much attention to the tool because I thought it would be more important to look at the pipeline issue… for the tool to return the collection of the coverage metric correctly”.
@@
In the last project, Organization C, there was the lowest access to the monitoring tool, with only 7 accesses to the dashboard by one participant. This could be explained by the company’s cultural resistance to CI adoption. In addition, the project had a holiday break at the end of the year, which is reflected in the access curve dropping and then rising again when organizational activities resumed. Another observation is that only one participant accessed the monitoring tool during the case study, which could potentially hinder a cultural shift regarding CI within the company.
@@
The registration of accesses to the dashboard allows us to conclude that Organizations A and B, which defined themselves as having an INTERMEDIATE CI maturity level, showed greater interest in accessing the tool. Organization C, which has a BASIC CI maturity level, exhibited less interest in using the tool.
IV-E RQ5: How did CI practices evolve during use of our metrics suite tool?
@@
Figure 3: (A) Organization A CI Practices Evolution during the case study. (B) Organization B CI Practices Evolution during the case study. (C) Organization C CI Practices Evolution during the case study.
@@
In Organization A, since the initiation of our monitoring efforts, all CI practices have demonstrated considerable improvement. Notably, the “Time to Fix a Broken Build” experienced a significant reduction, indicating enhanced efficiency in promptly addressing issues. As a consequence, the “Build Health” has increased significantly. The improvement in test coverage is equally relevant, and interviews revealed that monitoring practices played a crucial role in driving this positive trend. Participant P1 said: “I think the evaluation is good for us to have more detailed information, so we can make changes to the project, where necessary. We already knew about the coverage problem from the beginning, but it became very evident. With this information, we began to act to resolve this problem.”. The “Commits per Day” metric dropped in the middle of the case study, followed by a strong recovery. We associate this with the departure of a team member and, subsequently, the entry of new members into the development team. The same team change may have influenced the “Comments per Change” metric, which increased during the case study.
@@
Organization B presented a very unstable CI environment due to the complex integration with other applications, which was reflected in the CI metrics. Code coverage mostly remained zero despite the project having automated tests. Build time fluctuated greatly between normal weeks those with extremely high peaks (exceeding 80 hours). Participant P3 reflected on the issues with build time and coverage: “Yes, we have brought attention to these aspects in terms of build duration and code coverage. We are trying to understand this aspect in meetings, until this morning, the project managers asked me why the tests are taking so long”. This was a problem they were unable to resolve during the case study. This instability was also reflected in the high variation of “Build Health” and “Time to Fix a Broken Build” metrics, as Participant P2 declares: “Last week we had some problems with infrastructure, we ended up having to do some configurations and such. Because of it, the build health metric started to have a lower value”. The “Commits per Day” metric tended to decrease in the last four weeks of the case study compared to the first four weeks. This trend can be attributed to the conclusion of the development cycle, during which the project shifted its focus towards documentation, as reinforced by participant P4: “… as everyone is focused on documentation, there shouldn’t be any more commits, except something from the README, something very specific like that.”
@@
In Organization C, the project initially demonstrated good values for CI practices, which was surprising considering the self-declared BASIC level of CI maturity in the initial survey. However, there was a significant degradation in all metrics during the end-of-year holidays break (indicated by the red dotted circle). Despite this, the project experienced a quick recovery afterward in most metrics, such as: “Build Health”, “Build Activity” and “Time to Fix a Broken Build”. There is evidence that CI monitoring contributed to this recovery, as Participant P5 explained:: “There was also a period at the end of last year of the build being generated in a broken way, and as soon as we came back, one of the first things we did was fix it… We had already observed this in Gitlab, but the monitoring tool also contributed, in a way, complemented this need to fix it”
@@
Organization A experienced a notable improvement in CI metrics throughout the case study, particularly in coverage metric, where reports indicate that monitoring played a crucial role in identifying and addressing areas needing improvement. In contrast, Organization B faced challenges due to the instability of its CI environment, which impaired significant improvements in the project’s CI levels. Finally, in Organization C, the end-of-year holiday break caused a temporary decline in CI metrics, which were quickly improved with the help of monitoring.
IV-F Thematic Analysis
To conclude our case study, we performed a thematic analysis of the participants’ interviews. Figure 4 presents the main themes that emerged from the interviews about monitoring CI practices. The themes summarize ideas expressed by participants, and the number associated with each theme indicates how frequently it was mentioned. Green nodes represent positive aspects of monitoring CI practices, while red nodes indicate negative aspects highlighted by the interviewees.
Figure 4: Interviews Themes
The primary positive theme that emerged when monitoring CI practices was the “Desire for Improvements” (28 mentions), which illustrates the participant’s motivation to enhance certain CI practices in the project (i.e., a desire to improve specific practices in the future). For instance, participant P1 expressed this sentiment by stating: “We had already been doing this testing issue before, but it certainly reinforced the need to do it as quickly as possible.”. The second most cited theme was “Highlights the Problem” (23 mentions), that the monitoring tool helped the development team discover previously unnoticed issues. Participant P2 articulated this experience, saying:: “In fact, I found it strange that the coverage dropped, I thought it had already been resolved, there is something there that is being executed and resetting the data.”. The third most cited theme was “Ease of Monitoring” (18 mentions), which reflects how easily developers were able to monitor CI practices using our tool. Participant P4 commented: “No, no, it’s not expensive. It was very intuitive”. Finally, “Reminders and Motivators” (16 mentions) as another relevant theme, as the tool constantly highlighted issues to developers, motivating them to improve the CI process. As participant P3 remarked: “… but I think that seeing a very prominent percentage gets attention and helps convince other developers to write tests”.
Monitoring CI practices also provides an “Overview of CI State” for the project (12 mentions), enabling developers to view all CI-related aspects in one place, as participant P2 described: “… Precisely the tool summarizes several things that perhaps we need to use several tools to get to this information.”. Additionally, monitoring fosters “Encouraging Communication” among team members (12 mentions), as mentioned by participant P5: “I think it’s interesting to know what the practices are like. Just the fact that we talked somewhat superficially about commits would already be a benefit.”. Moreover, it promotes “CI Improvements” (9 mentions), as participant P1 emphasized: “I am more focused on developing the features that are missing and working on test coverage, which has the most impact for us. I just uploaded some testing issues that will improve the coverage”. Other themes were mentioned to a lesser extent, such as: “Impact on Project Quality” (4 mentions), “Encouraging Colaboration” (1 mention),“Gave a Direction” (1 mention), and “Encouraging Changes” (1 mention).
The primary negative theme in monitoring CI practices is “Challenges in Monitoring”. This emerged from difficulties using the tool or suggestions for improvement indicated by participants during the case study. Most of these issues were related to integrating the tool with other CI tools already used by the team, aiming to facilitate the CI practices monitoring. For example, participant P3 said: “If there was something like a link in the Gitlab interface as Sonarqube has. I could click and be redirected to the tool … ”, participant P5 commented: “I don’t know if the tool already has some type of integration with communication tools, could be a integration with the internal chat that we use for communication … ”. Integration with other tools is beyond the initial scope of this study, but it reflects the need for CI tools like GitLab to incorporate monitoring of CI practices directly into their pipelines.
The “Time Management Challenges” theme is mainly related to the difficulty teams face in implementing improvements to CI practices due to a lack of dedicated time. This is a common complaint, as developers usually have other responsibilities and cannot focus exclusively on improving CI practices for their projects. Overall, the monitoring tool proved to be easy and intuitive. Monitoring CI practices was not considered burdensome by participants. Participant P4 commented: “I think the tool is very intuitive, very didactic …”. Participant P1 said: “It has no major impact on configuration, doesn’t generate extra work, and adds significant value to this aspect of CI practices.”. Participant P5 appreciated the alerts sent by the tool: “I noticed that you can receive it by email. I found the feedback from the emails interesting.”
V Discussion
@@
In this section, we discuss the implications of our study. During the case study, it was observed that the automation of the build process and the execution of automated tests were the primary factors considered when defining the CI maturity level of projects. However, once CI monitoring was introduced, developers began to emphasize the monitoring process. Monitoring highlighted problems, served as a reminder, and motivated the team to enhance the CI process. They clearly perceived that such improvements could impact overall project quality.
@@
Combining CI practices into a single tool provides an overview of the CI process. It would be beneficial for tools like GitLab to incorporate CI practice monitoring. Such integration would assist in assessing CI maturity. The call for GitLab and similar platforms to support CI monitoring reflects a desire for more seamless and integrated solutions. Direct integration of CI monitoring into popular development platforms streamlines workflows, supports efficient collaboration, and provides real-time insights into the health and effectiveness of CI processes. This approach aligns with the broader industry shift towards DevOps practices, where collaboration between development and operations is tightly integrated.
@@
Implications for researchers. This case study showed that there are still gaps in monitoring development processes. Although CI is already a widespread practice, few studies seek to evaluate its efficient application. There is a field of study exploring new CI practices not evaluated in this study, as well as broader practices such as DevOps. We could also argue that other software engineering practices, in addition to CI, could potentially benefit from monitoring activities. For example, it would be interesting for future studies to monitor security-related metrics and present the results on an organized dashboard.
@@
Implications for practitioners. This study suggests that developers have an interest in the monitoring of CI practices and derive significant benefits from their adoption. Therefore, commercial CI services like GitHub, GitLab, or Jenkins would greatly benefit from incorporating dashboards to monitor development practices such as CI. This addition would add substantial value to these tools and contribute to improving the quality of projects utilizing them.
VI Threats to Validity
In this section, we discuss the limitations of our study and how we attempted to mitigate them.
@@ VI Threats to Validity
Scope: While our case study has provided valuable insights into Continuous Integration practices, its scope was limited to three software development organizations in our immediate geographic area, where we had physical access and permission to install the monitoring suite and collect data. The projects were represented by five participants. As a result, the generalization of our findings is constrained to the project contexts, and the participants may not have a complete understanding of CI across the entire project. We tried to select participants who had knowledge about the project and its CI infrastructure as a way to mitigate this limitation.
@@
Period of Case Study: The period during which the study was conducted — eight weeks — may not have been long enough to observe the full effects of any actions taken by the team after becoming aware of the project’s CI maturity. Furthermore, since we were immersed in the project environment, no developer was dedicated exclusively to the case study; they continued with their daily tasks and may not have had time to implement changes in the analyzed project, despite being alerted by the monitoring of CI practices
@@
The case study was carried out from November 2023 until February 2024, coinciding with the holiday season. Therefore, some projects were at the end of the development cycle, with lower demand; final tests or project documentation were carried out, which had a slight impact on some metrics during this period. A break in interviews was held between December 22, 2023, and January 19, 2024, to mitigate this issue.
Social Desirability Bias: To mitigate social desirability bias, we ensured a private and neutral environment for the interviews and clarified their purpose by emphasizing that all data collected would remain anonymous and be used solely for research purposes.
@@
Changes in the Project’s Team: Organization A experienced changes in its development team during the case study, which may have influenced the evolution of CI practice values reported in RQ5.
Insufficient Statistical Evidence: Although we observed an improvement in the metrics’ values at the end of the case study, future replication studies could help us collect enough data to confirm these hypotheses in a statistical manner.
@@
VII Conclusion
This study presents a case study of three software projects to explore the benefits of monitoring Continuous Integration (CI) practices in real-world settings — a topic often overlooked in both research and industry. To address this gap, the authors developed and deployed a minimally intrusive monitoring suite that allowed teams to monitor their CI practices. Results showed that CI monitoring is valuable for identifying issues, maintaining awareness of CI status, and motivating improvements. Developers expressed a strong interest in integrating monitoring dashboards with tools like GitLab, GitHub, and Jenkins to further streamline their workflows.
Future work should expand CI monitoring to encompass additional DevOps practices, including DORA metrics, to provide a more comprehensive view of the software delivery lifecycle. Integrating deployment and infrastructure metrics would align the monitoring framework with industry best practices and enable a more robust assessment of DevOps effectiveness. A possible direction is the development of a GitHub Action plugin based on the CI monitoring approach, which could be published on the GitHub Marketplace to encourage broader adoption.
VIII Acknowledgments
This work is partially supported by INES (www.ines.org.br), CNPq grant 465614/2014-0, CAPES grant 88887.136410/2017-00, and FACEPE grants APQ-0399-1.03/17 and PRONEX APQ/0388-1.03/14.